# Demystifying Network Foundation Models

### v2

@@ Abstract
This work presents a systematic investigation into the latent knowledge encoded within Network Foundation Models (NFMs) that focuses on hidden representations analysis rather than pure downstream task performance. Different from existing efforts, we analyze the models through a three-part evaluation: Embedding Geometry Analysis to assess representation space utilization, Metric Alignment Assessment to measure correspondence with domain-expert features, and Causal Sensitivity Testing to evaluate robustness to protocol perturbations. Using five diverse network datasets spanning controlled and real-world environments, we evaluate four state-of-the-art NFMs, revealing that they all exhibit significant anisotropy, inconsistent feature sensitivity patterns, an inability to separate the high-level context, payload dependency, and other properties. Our work identifies numerous limitations across all models and demonstrates that addressing them can significantly improve model performance (by up to +0.35 $F_{1}$ score without architectural changes).
@@ 1 Introduction
Network foundation models as a potential solution. As a potential answer to these challenges, network foundation models (NFMs) have been gaining traction in the networking community [59, 23, 32, 53, 41, 52]. Similar to foundation models in other application domains, these models incorporate additional self-supervised pretraining phases that utilize unlabeled data to learn critical spatial, temporal, and causal relationships. The adoption of NFMs promises to alleviate the difficulties caused by the growing complexity of modern network infrastructures and is viewed as an important step towards realizing the vision of self-driving networks.
Diverse architectures complicate evaluation. While aiming toward the similar goal of utilizing unlabeled data, design architectures and pretraining tasks vary significantly between existing NFMs, including masked token prediction purely on raw network payload [32] or packet header bytes [37, 52, 41], flow statistics calculation [23], patched image reconstruction [59, 53], modifying foundation models developed for other domains and applying them to the networking [51, 27]. Such diversity obscures the influence of the pretraining phase on model performance, making it difficult to understand what knowledge the model gains during pretraining. Historically, the community has relied on the performance of such models on fine-tuning tasks [40], which utilize small (compared to the pretraining phase) labeled datasets as quality indicators and comparison metrics between models. However, this approach poses challenges for understanding foundation models’ design choices, pretraining tasks, chosen pretraining datasets, and overall knowledge gained during the critical pretraining phase.
@@
Key findings. Our empirical results demonstrate (§4.3) that embeddings from publicly available pretrained models exhibit significant anisotropy (mean cosine similarity = $0.86\pm 0.09$), and that addressing this issue leads to model performance improvement (up to 0.35 $F_{1}$ score). We notice (§4.4) significant alignment of models’ embeddings with packet lengths, time-based features, packet flags, and even payload information (§4.5) despite its frequent encryption or absence in production environments, and show a direct correlation between anisotropy and high-level context in the models.
Our work shifts the focus from downstream task performance to intrinsic representational properties and sets the stage for further explorations into developing next-generation NFMs that could facilitate the creation of performant, generalizable, and robust ML-based solutions for disparate learning problems – a critical step toward realizing the ambitious goal of fully self-driving computer networks. Our fully reproducible code is available at https://github.com/maybe-hello-world/demystifying-networks.
@@ 2.2 Hidden context in networking
One key learning challenge in the networking domain is that network traffic is influenced by various hidden context $C$, which is not explicitly expressed in the network traffic traces but significantly influences it. This context consists of: various network conditions, which include partially observable characteristics of the network (throughput, latency), traffic shaping policies, network congestion situation, and influence of other network traffic passing through the same interfaces; application behavior which includes participants and their communication rules, such as communication protocols and congestion control algorithms, adaptive bitrate algorithms, application specific communication pattern (prevalence of download traffic over upload), inherent burstiness of the traffic; and others.
@@ 3 Intrinsic Evaluation Framework
Network foundation models (NFMs) pretrained on raw packet traces inherently encode an internal latent space reflecting various network and application characteristics (described in subsection 2.2) captured from the utilized training data. To gain insights into these characteristics, we explore embeddings generated from diverse network data by frozen NFMs.
Using the calculated embeddings, we aim to: quantify how effectively models distribute representations across their available dimensions (§3.1), demonstrating if they capture both shared protocol patterns and distinctive flow characteristics; measure the correspondence between learned embeddings and established network metrics (§3.2), revealing whether NFMs inherently calculate the same statistical features that domain experts have engineered over decades; and evaluate how embeddings respond to controlled perturbations at both protocol and contextual levels (§3.3), verifying that models learn meaningful causal dependencies. For each technique, we provide a claim (what NFMs should be able to do), rationale behind this claim, and our methodology to evaluate NFMs for these claims.
@@ 3.1 Embedding Geometry Analysis: quantifying representation space utilization
Rationale. Effective network modeling requires capturing both global protocol patterns and flow-specific variations. A well-distributed embedding geometry with measured anisotropy indicates [21] that the model distinguishes between flows along multiple meaningful dimensions. Such embeddings encode critical network characteristics including spatial phases (handshake vs. data vs. teardown), temporal patterns (short vs. long inter-arrival times), and causal reactions (packet loss vs. delays). Conversely, embeddings that cluster tightly with high cosine similarity suggest the model compresses diverse flows into nearly identical representations (similar to Li et al. [31], Su et al. [48]), overlooking subtle per-flow variations essential for robust network reasoning and downstream task performance.
@@ 3.2 Metric Alignment Assessment: measuring correspondence with domain-expert features
Methodology. We select a set of established network metrics calculated by CICFlowMeter [30], which have been extensively validated for traffic classification in prior works [2]. Let $\{x_{j}\}_{j=1}^{n}$ be our set of flows and let $m_{i}(x_{j})$ denote the value of the $i$th CICFlowMeter metric for flow $x_{j}$. We extract the embedding $h_{j}\in\mathbb{R}^{d}$ from the NFM’s final encoder layer and assemble $H=[h_{1},\ldots,h_{n}]$.
For each metric $m_{i}$, we compute the similarity index $\rho(m_{i},h_{j})$ between the metric and the model’s embedding representation across all traffic flows in the dataset using Centered Kernel Alignment (CKA) [29]. Unlike cosine similarity, CKA allows for calculating the similarity index between representations of different dimensionalities and is invariant to orthogonal transformations and isotropic scaling, allowing for effective capture of both linear and non-linear relationships between representations. We report the average CKA across all metrics. CKA values near 1 demonstrate that semantic probes successfully extract network-level metrics from embeddings, validating the model’s implicit computation of these key features.
@@ 3.3 Causal Sensitivity Testing: interventional analysis of protocol and context dependencies
A significant decrease in the similarity score after perturbation indicates that the model’s representation exhibits sensitivity to the perturbed features. Large changes for protocol-relevant features demonstrate that these features causally influence the model’s internal representations. This finding reveals that the model implicitly learns to encode protocol-relevant features, aligning with domain knowledge about network protocol behavior.
Methodology: extraction of exogenous network context. First, we use a network emulator to generate network traffic with distinct high‑level contexts ($c_{j}\in\{1,\dots,C\}$). Each context has multiple possible values, such as application type (video streaming vs. conferencing), congestion control algorithm (CUBIC vs. BBR), queue management policies (pFIFO vs. CoDEL), and cross traffic characteristics at the bottleneck link (more bursty and less intense vs. more intense and less bursty). For any given NFM, we extract embeddings $(h_{j})$ from the last encoder layer that represent the encoded state of $j^{th}$ flow. We compute these embeddings across different combinations of high-level contexts.
@@ 4.1 Network Foundation Models
For each model, we used publicly available pretrained checkpoints provided by the authors without modifications, maintaining each model’s original data processing pipeline to ensure fair comparison. Our selection includes YaTC [59], ET-BERT [32], netFound [23], and NetMamba [53]. The reader can refer to Appendix B for additional information about their architectures.
@@ 4.2 Datasets
Our benchmarking requires diverse network traffic datasets to ensure comprehensive evaluation across varying network conditions, traffic types, and operational settings. We selected five datasets spanning both controlled environments (Android Crossmarket [42], CIC-IDS2017 [46], CIC-APT-IIoT24 [22]) and real-world deployments (CAIDA [1], MAWI [13]). For preprocessing, we standardized all datasets by extracting uniform flow records using the data processing scripts provided by the corresponding model authors. The reader can refer to Appendix B for additional details.
@@ 4.3 Embedding Geometry Analysis
Table 1: Mean cosine similarity between embeddings inside the dataset and Mean Cosine Contribution (MCC) of the top dimension of the embeddings towards the average cosine similarity.
@@
Table 1 presents results from applying our embedding geometry analysis. The reader can refer to Appendix C for additional results.
Models are not consistent in the anisotropy scores. Our analysis reveals significant variability in how different NFMs utilize their representation space. netFound demonstrates more variability in space utilization (with anisotropy score varying between 0.69 and 0.94 for different datasets) compared to autoencoder-based models YaTC and NetMamba (which have more consistent scores of 0.85 and 0.99, respectively). ET-BERT occupies an intermediate position (0.71-0.88).
MCC analysis reveals distinct failure modes. While anisotropy scores identify models with concentrated representations, our Mean Cosine Contribution analysis further distinguishes between qualitatively different representational problems. NetMamba’s uniform MCC values (around 0.02) indicate evenly distributed but collectively collapsed representations, while YaTC’s uneven distribution (MCC 0.19-0.27) reveals problematic dimensional dominance where a single dimension (129) captures disproportionate variance.
Geometric properties predict environmental responses. Embedding geometry analysis systematically exposes how different architectures respond to dataset variations. The observation that most models produce higher cosine similarity (more collapsed representations) on exogenous datasets — with ET-BERT uniquely showing the opposite pattern — provides predictive insight into how these models might generalize to deployment environments with different traffic distributions.
@@
Anisotropy directly impacts performance. While anisotropy provides a measure of embedding distribution, it does not directly capture how this geometry affects downstream performance [16]. Therefore, we complement our anisotropy analysis with a practical measure: isotropification gain. This metric quantifies the potential performance improvement when correcting for suboptimal embedding distributions. We apply a deterministic decorrelation transformation [24] to the frozen raw representations and retrain only a linear classifier for the downstream task. We define the isotropification gain as $\Delta F_{1}=F_{1}^{\text{decorrelated}}-F_{1}^{\text{raw}}$. We also explore additional popular techniques, such as batch normalization and whitening, in Appendix D.
To validate that geometric properties directly affect model utility, we identified NetMamba as an ideal candidate for isotropification based on its high anisotropy. Table 2 shows the resulting $F_{1}$ score improvements (+0.03 to +0.35) after applying decorrelation transformations. This confirms that anisotropy is not merely a descriptive metric but can predict potential performance gains from representation enhancement techniques for foundational models in networking domain.
@@ 4.4 Metric Alignment Assessment
Table 3: CKA similarity index between CICFlowMeter white-box features and model embeddings.
@@
Table 3 presents the averaged Centered Kernel Alignment (CKA) similarity index between model embeddings and CICFlowMeter features across all datasets, quantifying the degree to which NFMs implicitly encode established network metrics without explicit supervision. The reader can refer to Appendix E for complete values for each feature.
Architectural design determines metric alignment. Our analysis reveals substantial variation in how different architectures encode domain-expert features. netFound consistently demonstrates the highest alignment (average CKA: 0.143), suggesting its multi-input approach and packet burst representation effectively capture statistical properties identified by domain experts. In contrast, ET-BERT’s payload-only approach shows minimal alignment (average CKA: 0.029), indicating traditional network metrics cannot be reliably extracted from payload representations alone.
Figure 1: CDF of CKA similarity among different model embeddings and CICFlowMeter features.
Feature diversity enhances metric coverage. Figure 1 visualizes the CDF of the CKA similarity index, revealing that models with diverse input features show more uniform alignment distribution. Models processing packet headers, payload, and flow metadata (netFound) capture a broader spectrum of expert-designed metrics compared to specialized architectures (ET-BERT, YaTC, NetMamba), indicating that architectural diversity directly impacts the comprehensiveness of implicit feature computation.
Environmental context affects metric relevance. All models demonstrate significantly lower CKA similarity on exogenous datasets (CAIDA: 0.032, MAWI: 0.055) compared to controlled environments (CIC-APT-IIoT24: 0.131, CIC-IDS2017: 0.091). This consistent pattern suggests existing NFMs struggle to encode established network metrics in production environments.
@@
Temporal-spatial features dominate alignment patterns. Table 4 reveals that all models demonstrate the highest alignment with features related to packet length, segment size, and flow burstiness. netFound shows particularly strong alignment with forward window bytes (CKA: 0.385) and packet length (CKA: 0.372), while YaTC prioritizes forward segment size (CKA: 0.423). Even ET-BERT, despite overall low alignment, shows relative preference for SYN flag counts (CKA: 0.122), suggesting all models discover the importance of initial connection behavior and data transfer volumes.

### v3

@@ Abstract
This work presents a systematic investigation into the latent knowledge encoded within Network Foundation Models (NFMs). Different from existing efforts, we focus on hidden representations analysis rather than pure downstream task performance and analyze NFMs through a three-part evaluation: Embedding Geometry Analysis to assess representation space utilization, Metric Alignment Assessment to measure correspondence with domain-expert features, and Causal Sensitivity Testing to evaluate robustness to protocol perturbations. Using five diverse network datasets spanning controlled and real-world environments, we evaluate four state-of-the-art NFMs, revealing that they all exhibit significant anisotropy, inconsistent feature sensitivity patterns, an inability to separate the high-level context, payload dependency, and other properties. Our work identifies numerous limitations across all models and demonstrates that addressing them can significantly improve model performance (up to 0.35 increase in $F_{1}$ scores without architectural changes).
@@ 1 Introduction
Network foundation models as a potential solution. As a potential answer to these challenges, network foundation models (NFMs) have been gaining traction in the networking community [59, 23, 32, 53, 41, 52]. Similar to foundation models in other application domains, these models incorporate additional self-supervised pretraining phases that utilize unlabeled data to learn critical spatial, temporal, and causal relationships. NFMs represent a paradigm shift in network analysis, moving from the development of task-specific learning models to the use of general-purpose pre-trained representations. Significant aspects of this shift include (i) generalizability: NFMs learn representations that transfer across multiple network analysis tasks; (ii) scale: NFMs process raw packet data and require no domain expert-based feature engineering; and (iii) impact: NFMs are expected to be increasingly deployed in production network systems for security, optimization, and monitoring. The adoption of NFMs promises to alleviate the difficulties caused by the growing complexity of modern network infrastructures and is viewed as an important step towards realizing the vision of self-driving networks.
Diverse architectures complicate evaluation. While aiming toward the similar goal of utilizing unlabeled data, design architectures and pretraining tasks vary significantly between existing NFMs, including masked token prediction purely on raw network payload [32] or packet header bytes [37, 52, 41]; flow statistics calculation [23]; patched image reconstruction [59, 53]; and modifying foundation models developed for other domains and applying them to the networking domain [51, 27]. Resulting largely from the variability in network data preprocessing (see Appendix A for additional details), such diversity obscures the influence of the pretraining phase on model performance, making it difficult to understand what knowledge the model gains during pretraining. Historically, the community has relied on the performance of such models on fine-tuning tasks [40], which utilize small (compared to the pretraining phase) labeled datasets as quality indicators and comparison metrics between models. However, this approach poses challenges for understanding foundation models’ design choices, pretraining tasks, chosen pretraining datasets, and overall knowledge gained during the critical pretraining phase.
@@
Key findings. Our empirical results demonstrate (§4.3) that embeddings from publicly available pretrained models exhibit significant anisotropy (mean cosine similarity = $0.86\pm 0.09$) and that addressing this issue leads to model performance improvement (up to 0.35 increase in $F_{1}$ scores). We notice (§4.4) significant alignment of models’ embeddings with packet lengths, time-based features, packet flags, and even payload information (§4.5) despite its frequent encryption or absence in production environments, and show a direct correlation between anisotropy and high-level context in the models.
Our work shifts the focus from downstream task performance to intrinsic representational properties and sets the stage for further explorations into developing next-generation NFMs that can be expected to facilitate the creation of performant, generalizable, and robust ML-based solutions for disparate learning problems – a critical step toward realizing the ambitious goal of fully self-driving computer networks. Our fully reproducible code is available at https://github.com/maybe-hello-world/demystifying-networks.
@@ 2.2 Hidden context in networking
One key learning challenge in the networking domain is that network traffic is influenced by various hidden context $C$, which is not explicitly expressed in the network traffic traces but significantly influences it. This context consists of: various network conditions, which include partially observable characteristics of the network (throughput, latency), traffic shaping policies, network congestion situation, and influence of other network traffic passing through the same interfaces; application behavior, which includes participants and their communication rules, such as communication protocols and congestion control algorithms, adaptive bitrate algorithms, application specific communication pattern (prevalence of download traffic over upload), inherent burstiness of the traffic; and others.
@@
2.3 Necessity of intrinsic evaluation
Since traditional evaluation efforts rely exclusively on downstream task performance, they suffer from critical shortcomings such as (i) performance scores provide no insight into what the models actually learn about network conditions or application behaviors, (ii) high performance for one task does not guarantee generalizability to new network domains, (iii) models can achieve good performance through dataset-specific shortcuts rather than genuine network understanding, and (iv) practitioners have no guidance for choosing models or identifying failure modes. Addressing these shortcomings calls for creating an intrinsic evaluation framework that can reveal what the models actually learn about network traffic, independent of specific tasks.
@@ 3 Intrinsic Evaluation Framework
Intrinsic Evaluation FrameworkOriginalNetwork TrafficPerturbedNetwork TrafficWhite-boxFeature ExtractorsNetworkFoundation ModelWhite-boxFeaturesOriginalEmbeddingsPerturbedEmbeddingsII: Metric AlignmentAssessmentI: EmbeddingGeometry AnalysisIII: CausalSensitivity Testing
Figure 1: Visual overview of the proposed framework. Color indicates different stages of the analysis and dashed lines and boxes denote the perturbed network traffic path.
NFMs pretrained on raw packet traces inherently encode an internal latent space reflecting various network and application characteristics captured from the utilized training data. To gain insights into these characteristics, we explore embeddings generated from diverse network data by frozen NFMs.
Using the calculated embeddings, we aim to: quantify how effectively models distribute representations across their available dimensions (§3.1), demonstrating if they capture both shared protocol patterns and distinctive flow characteristics; measure the correspondence between learned embeddings and established network metrics (§3.2), revealing whether NFMs inherently calculate the same statistical features that domain experts have engineered over decades; and evaluate how embeddings respond to controlled perturbations at both protocol and contextual levels (§3.3), verifying that models learn meaningful causal dependencies. For each technique, we provide a claim (what NFMs should be able to do), the rationale behind this claim, and our methodology to evaluate NFMs for the given claim. A visual overview of the Intrinsic Evaluation Framework is presented in Figure 1.
Applicability. In theory, our intrinsic evaluation framework can be used to evaluate any model that learns latent space representations, but it is most valuable for models that involve task-agnostic pre-training. While traditional approaches (e.g., neural networks trained using supervised learning techniques) can be evaluated using our metric alignment component, they often lack the explicitly extracted latent representations that make geometric analysis and causal sensitivity testing meaningful.
Importantly, because of the data representation challenges posed by network traffic (see Appendix A), how a model’s input data is being preprocessed is intrinsically entangled with what model architecture is selected. Since both aspects reflect the explicit decisions that the original model’s developers made, we are unable to evaluate them individually.
@@ 3.1 Embedding Geometry Analysis: quantifying representation space utilization
Rationale. Effective network modeling requires capturing both global protocol patterns and flow-specific variations. A well-distributed embedding geometry with measured anisotropy indicates [21] that the model distinguishes between flows along multiple meaningful dimensions. Such embeddings encode critical network characteristics including spatial (handshake vs. data exchange vs. teardown) or temporal patterns (short vs. long inter-arrival times) and causal reactions (packet loss vs. delays). Conversely, embeddings that cluster tightly with high cosine similarity suggest the model compresses diverse flows into nearly identical representations (similar to Li et al. [31], Su et al. [48]), overlooking subtle per-flow variations essential for robust network reasoning and downstream task performance.
@@ 3.2 Metric Alignment Assessment: measuring correspondence with domain-expert features
Methodology. We select a set of established network metrics calculated by CICFlowMeter [30], which have been extensively validated for traffic classification in prior works [2]. Let $\{x_{j}\}_{j=1}^{n}$ be our set of flows and let $m_{i}(x_{j})$ denote the value of the $i$th CICFlowMeter metric for flow $x_{j}$. We extract the embedding $h_{j}\in\mathbb{R}^{d}$ from the NFM’s final encoder layer before any task-specific heads (ET-BERT, netFound) or decoder layers (YaTC, NetMamba) and assemble $H=[h_{1},\ldots,h_{n}]$.
For each metric $m_{i}$, we compute the similarity index $\rho(m_{i},h_{j})$ between the metric and the model’s embedding representation across all traffic flows in the dataset using Centered Kernel Alignment (CKA) [29]. Unlike cosine similarity, CKA allows for calculating the similarity index between representations of different dimensionalities and is invariant to orthogonal transformations and isotropic scaling, allowing for effective capture of both linear and non-linear relationships between representations. CKA values near 1 demonstrate that semantic probes successfully extract network-level metrics from embeddings, validating the model’s implicit computation of these key features.
@@ 3.3 Causal Sensitivity Testing: interventional analysis of protocol and context dependencies
A significant decrease in the similarity score after perturbation indicates that the model’s representation exhibits sensitivity to the perturbed features. Large changes for protocol-relevant features demonstrate that these features causally influence the model’s internal representations. Such findings reveal that the model implicitly learns to encode protocol-relevant features, aligning with domain knowledge about network protocol behavior.
Methodology: extraction of exogenous network context. First, we use a network emulator to generate network traffic with distinct high‑level contexts ($c_{j}\in\{1,\dots,C\}$). Each context has multiple possible values, such as application type (video streaming vs. conferencing), congestion control algorithm (CUBIC vs. BBR), queue management policies (pFIFO vs. CoDEL), and cross traffic characteristics at the bottleneck link (more bursty and less intense or vice versa). For any given NFM, we extract embeddings $(h_{j})$ from the last encoder layer that represent the encoded state of $j^{th}$ flow. We compute these embeddings across different combinations of high-level contexts.
@@ 4.1 Network Foundation Models
For each model, we used publicly available pretrained checkpoints provided by the authors without modifications, maintaining each model’s original data processing pipeline to ensure fair comparison. Our selection includes YaTC [59], ET-BERT [32], netFound [23], and NetMamba [53]. Additional information about their architectures can be found in Appendix B.
@@ 4.2 Datasets
Our benchmarking requires diverse network traffic datasets to ensure comprehensive evaluation across varying network conditions, traffic types, and operational settings. We selected five datasets spanning both controlled environments (Android Crossmarket [42], CIC-IDS2017 [46], CIC-APT-IIoT24 [22]) and real-world deployments (CAIDA [1], MAWI [13]). For preprocessing, we standardized all datasets by extracting uniform flow records using the data processing scripts provided by the authors of the considered models. We refer the reader to Appendix B for additional details.
@@ 4.3 Embedding Geometry Analysis
Table 1: Mean cosine similarity (cos) between embeddings and Mean Cosine Contribution (MCC) of the top dimension of the embeddings towards the average cosine similarity.
@@
Table 1 presents results from applying our embedding geometry analysis (additional results can be found in Appendix C).
Models are not consistent in the anisotropy scores. Our analysis reveals significant variability in how different NFMs utilize their representation space. netFound demonstrates more variability in space utilization (with anisotropy score varying between 0.69 and 0.94 for different datasets) compared to autoencoder-based models YaTC and NetMamba (which have more consistent scores of 0.85-0.88 and 0.92-0.99, respectively). ET-BERT occupies an intermediate position (0.71-0.88).
MCC analysis reveals distinct failure modes. While anisotropy scores identify models with concentrated representations, our Mean Cosine Contribution analysis further distinguishes between qualitatively different representational problems. NetMamba’s uniform MCC values (the highest being around 0.03) indicate semantically collapsed representations without any single dimension contributing significantly to the collapse, while YaTC’s uneven distribution (MCC 0.19-0.27) reveals problematic dimensional dominance where a single dimension captures disproportionate variance.
Geometric properties predict environmental responses. Embedding geometry analysis systematically exposes how different architectures respond to dataset variations. The observation that most models produce higher cosine similarity (more collapsed representations) on real-world datsets — with ET-BERT uniquely showing the opposite pattern — provides predictive insight into how these models might generalize to deployment environments with different traffic distributions.
@@
Anisotropy directly impacts performance. While anisotropy is a measure that quantifies embedding quality, it does not directly capture how this geometry affects downstream performance [16]. Therefore, we complement our anisotropy analysis by considering an additional metric: isotropification gain. This metric quantifies the potential performance improvement when correcting for suboptimal embedding distributions. We apply a deterministic decorrelation transformation [24] to the frozen raw representations and retrain only a linear classifier for the downstream task. We define the isotropification gain as $\Delta F_{1}=F_{1}^{\text{decorrelated}}-F_{1}^{\text{raw}}$. Additional popular techniques, such as batch normalization and whitening, are discussed in Appendix D.
To validate that embedding geometry directly affects model utility, we identified NetMamba as an ideal candidate for isotropification based on its high anisotropy. Table 2 shows the resulting $F_{1}$ score improvements (+0.03 to +0.35) after applying decorrelation transformations. Note that this improvement is calculated for a simplified classification head (linear classifier), and results for more complicated models can vary. However, these results suggest that anisotropy is not merely a descriptive metric but can predict potential performance gains that derive from applying representation enhancement techniques to NFMs.
@@ 4.4 Metric Alignment Assessment
Table 3: Averaged CKA similarity index between CICFlowMeter white-box features and model embeddings.
@@
Table 3 presents the averaged Centered Kernel Alignment (CKA) similarity index between model embeddings and CICFlowMeter features across all datasets, quantifying the degree to which NFMs implicitly encode established network metrics without explicit supervision. Appendix E provides a breakdown of this average similarity index into the similarity indices of the individual features.
Figure 2: CDF of CKA similarity among different model embeddings and CICFlowMeter features averaged across all five datasets.
Architectural design determines metric alignment. Our analysis reveals substantial variation in how different architectures encode domain-expert features. netFound consistently demonstrates the highest alignment (average CKA: 0.143), suggesting its multi-input approach and packet burst representation effectively capture statistical properties identified by domain experts. In contrast, ET-BERT’s payload-only approach shows minimal alignment (average CKA: 0.029), indicating traditional network metrics cannot be reliably extracted from payload representations alone. Since YaTC and NetMamba share similar architectural designs, they demonstrate similar CKA performance across datasets.
Feature diversity enhances metric coverage. Figure 2 shows the CDF of the CKA similarity index, revealing that models with diverse input features exhibit a more uniform alignment distribution compared to models with constrained input features (i.e., more slowly increasing CDF, resulting in more features having a high CKA similarity value). Models processing packet headers, payload, and flow metadata (netFound) capture a broader spectrum of expert-designed metrics compared to specialized architectures (ET-BERT, YaTC, NetMamba), indicating that architectural diversity directly impacts the comprehensiveness of implicit feature computation.
Environmental context affects metric relevance. All models demonstrate significantly lower CKA similarity on real-world datasets (CAIDA: 0.032, MAWI: 0.055) compared to datasets from controlled environments (CIC-APT-IIoT24: 0.131, CIC-IDS2017: 0.091). This consistent pattern suggests that existing NFMs struggle to encode established network metrics in production environments.
@@
Temporal-spatial features dominate alignment patterns. Table 4 shows that all models demonstrate the highest alignment with features related to packet length, segment size, and flow burstiness. For instance, netFound shows particularly strong alignment with features forward initial window bytes (CKA: 0.385) and packet length max (CKA: 0.372), while YaTC prioritizes the feature forward segment size min (CKA: 0.423). Even ET-BERT, despite overall low alignment, shows relative preference for the feature SYN flag count (CKA: 0.122). These observations suggest that all models discover the importance of learning initial connection behavior and data transfer volumes.